OpenStudy (anonymous):

Arectangle has a perimeter of 30m. The length is 10m. Solve the following equaion to find the width. 2w+2(10)=30

jimthompson5910 (jim_thompson5910):

2(10) = ???

OpenStudy (anonymous):

20

jimthompson5910 (jim_thompson5910):

so 2w+2(10)=30 2w+20=30

jimthompson5910 (jim_thompson5910):

then what

OpenStudy (anonymous):

does that mean w=10

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

Can you explain it I really dont get this

jimthompson5910 (jim_thompson5910):

2w+2(10)=30 2w+20=30 2w+20-20=30-20 2w=10 2w/2=10/2 w=5 So it's really w = 5